Lyn's father is 4 times as old as her. Five years ago the father was 7 times as old as lyn. How old is lyn now?
Answer by ikleyn(52864)

.
Let x be the Lyn's age.

Then the father's age is 4x.


Five years ago Lyn was (x-5) years old, while her father was (4x-5) years old.


The condition says

4x-5 = 7(x-5).


Simplify and solve


4x - 5 = 7x - 35

35 - 5 = 7x - 4x

30 = 3x

x = 30/3 = 10.


Answer.  Lyn is 10 years old now.

Solved.